# Jensen’s Functional Equation on Involution-Generated Groups: An ($\mathrm{SR}_{2}$) Criterion and Applications

### v1

@@ Abstract.
with the normalization $f(e)=0$. Building on techniques for the symmetric groups $S_{n}$, we isolate a structural criterion on GG—phrased purely in terms of involutions and square roots—under which every solution to (J1) must also satisfy (J2) and is automatically a group homomorphism. Our new criterion, denoted $(\mathrm{SR}_{2})$, implies that $S_{1}(G,H)=S_{1,2}(G,H)=\mathrm{Hom}(G,H)$, applies to many reflection–generated groups and, in particular, recovers the full solution on $S_{n}$. Furthermore, we give a transparent description of the solution space in terms of the abelianization $G/[G,G]$, and we treat dihedral groups $D_{m}$ in detail, separating the cases $m$ odd vs. even. The approach is independent of division by $2$ in $H$ and complements the classical complex-valued theory that reduces (J1) to functions on $G/[G,[G,G]]$.
@@ 1. Introduction
Throughout this paper, the group $G$ is written multiplicatively, while the abelian group $H$ is written additively. We also normalize our function by assuming that $f(e)=0.$
@@ 2. A structural square-root criterion
Let $G$ be a group generated by involutions and $H$ an abelian group. Assume $f:G\to H$ satisfies (J1) with $f(e)=0$. If $2f\equiv 0$ on $G,$ then:
@@
Applying (J1) to $(xy^{-1},z)$ and $(xz^{-1},y)$ and using the hypothesis $2f\equiv 0$ shows that $f(xy^{-1}z^{-1})=-f(xy^{-1}z)$ and $f(xz^{-1}y^{-1})=-f(xz^{-1}y)$. Hence the difference becomes
@@
Step B: reordering and absorption tools. By Theorem 2.7, $f(Zt^{2}j)=f(Zjt^{2})$ for all $Z,t$. Applying (J1) to $(x,y)=(Zt,t)$ and using Lemma 2.6 ($2f\equiv 0$ on words of involutions) yields the absorption identity

### v2

@@ Abstract.
with the normalization $f(e)=0$. Building on techniques for the symmetric groups $S_{n}$, we isolate a structural criterion on $G$, phrased purely in terms of involutions and square roots, under which every solution to (J1) must also satisfy (J2) and is automatically a group homomorphism. Our new criterion, denoted $(\mathrm{SR}_{2})$, implies that $S_{1}(G,H)=S_{1,2}(G,H)=\mathrm{Hom}(G,H)$, applies to many groups generated by reflections, and in particular recovers the full solution on $S_{n}$. Furthermore, we give a transparent description of the solution space in terms of the abelianization $G/[G,G]$, and we treat dihedral groups $D_{m}$ in detail, separating the cases $m$ odd vs. even. The approach is independent of division by $2$ in $H$ and complements the classical complex-valued theory that reduces (J1) to functions on $G/[G,[G,G]]$.
@@ 1. Introduction
Throughout this paper, the group $G$ is written multiplicatively, while the abelian group $H$ is written additively. We also normalize our function by assuming that $f(e)=0.$ Note that every solution of (J1) is of the form $f=f_{0}+c$ with $f_{0}(e)=0$ and $c\in H$ constant; thus there is no loss of generality in working with normalized solutions throughout.
@@ 2. A structural square-root criterion
Let $G$ be a group generated by involutions and $H$ an abelian group. Assume $f:G\to H$ satisfies (J1) with $f(e)=0$. Then:
@@
Applying (J1) to $(xy^{-1},z)$ and $(xz^{-1},y)$ and using Lemma 2.6 (which gives $2f(xy^{-1})=2f(xz^{-1})=0$) shows that
$f(xy^{-1}z^{-1})=-\,f(xy^{-1}z),\quad f(xz^{-1}y^{-1})=-\,f(xz^{-1}y).$
Hence the difference becomes
@@
Step B: reordering and absorption tools.
By Theorem 2.7, we may reorder the factors in the argument of $f$; in particular $f(Zt^{2}j)=f(Zjt^{2})$ for all $Z,t$.
Applying (J1) to $(x,y)=(Zt,t)$ and using Lemma 2.6 ($2f\equiv 0$ on words of involutions) yields the absorption identity